# Fokker-Planck approach for thermal fluctuations in antiferromagnetic systems

### v1

@@ IV.2 Thermal fluctuations in the resistance
We conclude with a discussion about the full resistance power spectrum. Fig. 4 a) shows the resistance power spectrum as a function of the temperature for different values of $\omega$ (color-coded) and the dimensionless parameters $\delta\lambda/\lambda$ (line-style-coded). We observe that, independently of the value of $\delta\lambda$, the contribution of $S_{\parallel}(\omega)$ dominates over $S_{\bot}(\omega)$, as is evident when comparing each curve with their corresponding in Fig. 2 a). Note in particular the maximum below $T_{\rm N}$ for $\omega=\lambda J/10$ and $\omega=\lambda J$. Switching on the cylindrical symmetry in the Gilbert damping factor, $\delta\lambda\neq 0$ slightly modifies the amplitude of the resistance power spectrum. Finally, we comment on Fig. 4 b), which shows the $\omega$-dependence of the resistance power spectrum for several temperatures (color-coded) and values of $\delta\lambda/\lambda$ (line-style-coded). In this log-log plot, the decreasing trend as $1/\omega^{2}$ (dotted gray line) for $\omega\gtrsim\lambda J$ is expected, since $S_{R}(\omega)$ is the sum of Lorentzian functions.
The behavior described by this phenomenological model qualitatively resembles that reported in ghosh2024electronic, where the authors observed an unusual feature in the noise resistance data obtained from measurements in a 2D AFM system, superimposed on the expected $1/f$ noise paladino20141; pellegrino_apl_2023; pellegrino_jstat_2019; pellegrino_commphys_2020; hardner_prb_1993. Specifically, below the Néel temperature, the resistance power spectrum exhibits a contribution that can be described by a Lorentzian function with a characteristic damping rate $\Gamma\gtrsim 10$$10^{2}~{\rm Hz}$, corresponding to the typical range of low-frequency resistance noise measurements, e.g. $\omega=10$ Hz. This feature is most pronounced slightly below the Néel temperature, resembling the red and green curves in Fig. 4 a).
@@ V Conclusions
We further applied this formalism to model spin-wave dynamics, finding a renormalization of the characteristic energies and the damping rates due to the thermal fluctuations. Moreover, we have defined a phenomenological model that describes the influence of magnetic fluctuations on the resistance noise of 2D antiferromagnetic semiconductors. We derived analytic expressions for the resistance power spectrum, showing that thermal fluctuations generate Lorentzian noise contributions with characteristic damping rates. This model exhibits key qualitative characteristics that resemble the measurements observed experimentally ghosh2024electronic, in particular, the occurrence of Lorentzian low-frequency resistance noise maximized near the AFM transition.
@@ Appendix A Kramers-Moyal expansion and the Fokker-Planck equation
Within this Appendix, we start from Eq. (7), which represents a Langevin equation. The corresponding stochastic process is Markovian, and its probability distribution function (PDF) satisfies a master equation, that we derive here in the Kramers–Moyal form kramers1940brownian; moyal1949quantum; garcia2007introduction. In general, the Kramers–Moyal expansion transforms a master equation into infinite-order differential equation. When truncated at the second-order, the resulting partial differential equation is known as the Fokker–Planck (FP) equation. In the present case, by employing (9) of the main-text, the truncation is exact garcia2007introduction, and the Kramers-Moyal coefficients are expressed as
@@
Following Ref. kamra_prb_2018, we introduce the inter-sublattice Gilbert damping terms via the diagonal matrix $\Lambda^{\prime}$.
@@
from a comparison with the local minimum of the free energy via the classical Hamiltonian, using the MF method goldenfeld_book, given by $\frac{B(\xi)}{\xi}=\frac{k_{\rm B}T}{J}$, one has
@@ Appendix B Equation of motion of correlation functions
where $\mathcal{P}_{0}({\bf\it\mathcal{S}})$ represents the probability distribution in the stationary regime likharev2019statistical. In the case with $N=1$ and $M=1$, one finds the correlation functions defined with Eq. (23) of the main-text.
@@
By using the FP equation on ${\mathcal{P}}({\bf\it\mathcal{S}},t|{\bf\it\mathcal{S}}^{\prime},0)$, the resulting equation of motion for the correlation function is obtained garcia2007introduction; garanin1997fokker. This equation has the same structure of the equation for the dynamics of the magnetization, with the precession term, the diffusion term, and finally the intra-sublattice and inter-sublattice terms

### v2

@@ IV.2 Thermal fluctuations in the resistance
We conclude with a discussion about the full resistance power spectrum. Fig. 5 a) shows the resistance power spectrum as a function of the temperature for different values of $\omega$ (color-coded) and the dimensionless parameters $\delta\lambda/\lambda$ (line-style-coded). We observe that, independently of the value of $\delta\lambda$, and even under the assumption that the resistance depends spherically symmetrically on the net field $\bm{h}$} ($\kappa_{R}^{\prime}=0$), the contribution of $S_{\parallel}(\omega)$ dominates over $S_{\bot}(\omega)$, as is evident when comparing each curve with their corresponding in Fig. 3 a). Note in particular the maximum below $T_{\rm N}$ for $\omega=\lambda J/10$ and $\omega=\lambda J$. Switching on the cylindrical symmetry in the Gilbert damping factor, $\delta\lambda\neq 0$ slightly modifies the amplitude of the resistance power spectrum. Finally, we comment on Fig. 5 b), which shows the $\omega$-dependence of the resistance power spectrum for several temperatures (color-coded) and values of $\delta\lambda/\lambda$ (line-style-coded). In this log-log plot, the decreasing trend as $1/\omega^{2}$ (dotted gray line) for $\omega\gtrsim\lambda J$ is expected, since $S_{R}(\omega)$ is the sum of Lorentzian functions.
The behavior described by this phenomenological model qualitatively resembles that reported in [15], where the authors observed an unusual feature in the noise resistance data obtained from measurements in a 2D AFM system, superimposed on the expected $1/f$ noise [38, 43, 41, 42, 20]. Specifically, below the Néel temperature, the resistance power spectrum exhibits a contribution that can be described by a Lorentzian function with a characteristic damping rate $\Gamma\gtrsim 10$$10^{2}~{\rm Hz}$, corresponding to the typical range of low-frequency resistance noise measurements, e.g. $\omega=10$ Hz. This feature is most pronounced slightly below the Néel temperature, resembling the red and green curves in Fig. 5 a).
@@ V Conclusions
We further applied this formalism to model spin-wave dynamics, finding a renormalization of the characteristic energies and the damping rates due to the thermal fluctuations. Moreover, we have defined a phenomenological model that describes the influence of magnetic fluctuations on the resistance noise of 2D antiferromagnetic semiconductors. We derived analytic expressions for the resistance power spectrum, showing that thermal fluctuations generate Lorentzian noise contributions with characteristic damping rates. This model exhibits key qualitative characteristics that resemble the measurements observed experimentally [15], in particular, the occurrence of Lorentzian low-frequency resistance noise maximized near the AFM transition.
@@ Appendix A Kramers-Moyal expansion and the Fokker-Planck equation
Within this Appendix, we start from Eq. (13), which represents a Langevin equation. The corresponding stochastic process is Markovian, and its probability distribution function (PDF) satisfies a master equation, that we derive here in the Kramers–Moyal form [27, 35, 12]. In general, the Kramers–Moyal expansion transforms a master equation into infinite-order differential equation. When truncated at the second-order, the resulting partial differential equation is known as the Fokker–Planck (FP) equation. In the present case, by employing (15) of the main-text, the truncation is exact [12], and the Kramers-Moyal coefficients are expressed as
@@
Following Refs. [23, 24], we introduce the inter-sublattice Gilbert damping terms via the diagonal matrix $\Lambda^{\prime}$.
@@
from a comparison with the local minimum of the free energy via the classical Hamiltonian, using the MF method [18], given by $\frac{B(\xi)}{\xi}=\frac{k_{\rm B}T}{J}$, one has
@@ Appendix B Equation of motion of correlation functions
where $\mathcal{P}_{0}({\bf\it\mathcal{S}})$ represents the probability distribution in the stationary regime [30]. In the case with $N=1$ and $M=1$, one finds the correlation functions defined with Eq. (30) of the main-text.
@@
By using the FP equation on ${\mathcal{P}}({\bf\it\mathcal{S}},t|{\bf\it\mathcal{S}}^{\prime},0)$, the resulting equation of motion for the correlation function is obtained [12, 11]. This equation has the same structure of the equation for the dynamics of the magnetization, with the precession term, the diffusion term, and finally the intra-sublattice and inter-sublattice terms